OpenStudy (anonymous):

at what distance above the surface of the earth is the acceleration due to gravity 0.980m/s² if the acceleration due to gravity at the surface has magnitude 9.80m/s²?

OpenStudy (anonymous):

30635

OpenStudy (ybarrap):

$$ \large{ F =mg= G\frac{m\times m_{earth}}{r^2}\\ g=G\frac{m_{earth}}{r^2}\ } $$ So @ 10% of g $$ \large{ 0.980=G\frac{m_{earth}}{r^2}\\ \implies r=\sqrt{\cfrac{G m_{earth}}{0.980}} } $$ Distance from the surface is $$ \large{ r-R_{earth} } $$ Where \(R_{earth}\) is the radius of the Earth. This turns out to be about 1800 kilometers or 1118 miles.